Below are hypothetical data from a titration experiment. Part A focuses on determining the molarity of
a sodium hydroxide solution. Part B uses the sodium hydroxide solution to determine the molarity, and
ultimately the mass percent, of acetic acid in an unknown vinegar sample. Vinegar is an aqueous
solution containing 4 to 6% by mass acetic acid. Please use the data provided below to perform the
calculations and answer the questions in this assignment.

Part A:

A 0.883 M standard HCl solution was used to standardize a solution of sodium hydroxide. Below are the
raw titration data for three trials with the HCl and sodium hydroxide.

Titration 1: A 5.00 mL aliquot of HCl solution was added to a flask. The flask was titrated with a sodium
hydroxide solution (in the buret). The initial reading of the buret was 1.20 mL. The final reading was
40.38 mL.

Titration 2: A 5.00 mL aliquot of HCl solution was added to a flask. The flask was titrated with a sodium
hydroxide solution (in the buret). The initial reading of the buret was 0.79 mL. The final reading was
40.10 mL.

Titration 3: A 5.00 mL aliquot of HCl solution was added to a flask. The flask was titrated with a sodium
hydroxide solution (in the buret). The initial reading of the buret was 1.00 mL. The final reading was
40.20 mL.

Part B:

Using the sodium hydroxide solution from Part A, an unknown sample of vinegar (#53) was titrated to
determine the molarity of acetic acid in the vinegar. Below are the raw titration data for three trials
with the sodium hydroxide and the vinegar samples.

Titration 1: A 5.00 mL aliquot of vinegar was added to a flask. The flask was titrated with a sodium
hydroxide solution (in the buret). The initial reading of the buret was 0.57 mL. The final reading was
35.05 mL.

Titration 2: A 5.00 mL aliquot of vinegar was added to a flask. The flask was titrated with a sodium
hydroxide solution (in the buret). The initial reading of the buret was 0.88 mL. The final reading was
34.35 mL.

Titration 3: A 5.00 mL aliquot of vinegar was added to a flask. The flask was titrated with a sodium
hydroxide solution (in the buret). The initial reading of the buret was 0.72 mL. The final reading was
35.14 mL.
Answer #1

For part A, we calculate the concentration of NaOH in each titration with the equation:

Cb = Ca * Va / Vb = 0.883 M * 5 mL / (40.38 - 1.20) mL = 0.1127 M

Trial 2: 0.1123 M

Trial 3: 0.1126 M

Average = 0.1125 M

Part B: The concentration of acetic acid is calculated:

Ca = Cb * Vb / Va = 0.1125 M * (35.05 - 0.57) mL / 5 mL = 0.7758 M

Trial 2: 0.7531 M

Trial 3: 0.7745 M

Average: 0.7678 M

The percentage of acetic acid in solution is calculated:

% AA = 0.7678 mol / L * (1 L / 1000 mL) * (60.05 g / 1 mol) * 100 = 4.61%

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.